# Thunderscapes: Simulating the Dynamics of Mesoscale Convective System

### v3

@@ 4.2. Electrification
Based on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair-weather electric field ($E$) induces opposite charges on precipitation particles. Collisions between particles, influenced by their velocities and directions, result in partial charge neutralization and a residual net charge. This theory underpins the following mathematical model for electrification, with charge density ($\rho$) defined as:
(16) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}v_{c}-q_{p}^{2}v_{p}\right)e^{-10q_{v}}$
where $k$ is a user-defined constant, $v_{c}$ and $v_{p}$ are their respective velocities of cloud-ice mixture and rain-snow mixture.
@@
The threshold electric field ($E_{tr}$) required for lightning initiation depends on the altitude ($\rho_{A}(h)$) as defined by (Marshall et al., 1995):
(17) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right)$
@@
Charge neutralization processes, occurring after a lightning discharge as described by (Barthe and Pinty, 2007), are governed by the following relationship:
(18) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}}\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}}\end{cases}$
@@
Here, $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density. The total neutralized charge density, accounting for the collective contributions of lightning growth points, is expressed as:
(19) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho$
@@
Finally, a suppression factor ($S$) is introduced to modulate lightning activity frequency as thunderstorms dissipate. This factor ensures consistent lightning activity during the lifecycle of thunderstorms:
(20) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}$
@@ 4.3. Atmospheric Background
(21) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}}$
@@
(22) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1}\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h\end{cases}$
@@
(23) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}}$
@@
(26) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}$
@@
4.4. Electrodynamics
The electrodynamic behavior of the atmosphere is governed by Maxwell’s equations(Ma et al., 1998), which define the relationships between the electric field $\mathbf{E}$, charge density $\rho$, and electric potential $V$. Gauss’s law describes how the divergence of $\mathbf{E}$ is proportional to $\rho$:
(28) $\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}}$
where $\epsilon_{0}$ is the permittivity of free space. The electric field is related to the potential $V$ through:
(29) $\mathbf{E}=-\nabla V$
indicating that $\mathbf{E}$ points in the direction of decreasing potential. Together, these equations describe the electrostatic field and potential in regions with charge density.
To analyze the electric potential in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics. This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
(30) $\nabla^{2}V=-4\pi\rho$
For simulating lightning discharges, the Dielectric Breakdown Model (DBM) is employed(Niemeyer et al., 1984). The probability of a discharge at a specific point $i$ depends on the electric potential $V_{i}$, expressed as:
(31) $p_{i}=\frac{\left(V_{i}\right)^{\eta}}{\sum_{j=1}^{n}\left(V_{j}\right)^{\eta}}$
Here, $\eta$ is a parameter controlling the spatial concentration of the discharge(Kim and Lin, 2007). Larger values of $\eta$ bias the discharge toward regions of higher potential, while smaller values allow for more diffusive propagation.
@@ 4.5. Fluid Dynamics
The motion of atmospheric fluids is described by the Navier-Stokes equations, which represent the conservation of momentum in a fluid medium (Wendt, 2008). These equations capture the effects of inertial forces, pressure gradients, viscosity, and external forces:
(32) $\frac{\partial(\rho\mathbf{u})}{\partial t}+\nabla\cdot(\rho\mathbf{u}\mathbf{u})=-\nabla p+\rho\left(\mu\nabla^{2}\mathbf{u}+\mathbf{b}+\mathbf{f}\right),\quad\nabla\cdot\mathbf{u}=0$
Here, $\rho$ is the fluid density, $\mathbf{u}$ the velocity field, $p$ the pressure, $\mu$ the dynamic viscosity, $\mathbf{b}$ the body force (e.g., gravity), and $\mathbf{f}$ external forces. The incompressibility condition $\nabla\cdot\mathbf{u}=0$ ensures mass conservation for incompressible flows.
@@ 5. ALGORITHMICS
Algorithm 1 Thunderstorm Development Algorithm.
1:Input: Current MCS state $(\theta,p,\rho,u,q_{v},q_{c},q_{p})$.
@@
4:     $\theta,p,\rho\leftarrow$ Update atmospheric background conditions Eqs.(22–23)
5:     $u\leftarrow$ Advect and diffuse velocity field
6:     $b\leftarrow$ Compute thermal buoyancy Eq.(27)
7:     $u\leftarrow u+b+f_{w}+f_{v}$triangleright Apply buoyancy, wind, vorticity confinement
@@
9:     $u\leftarrow$ Pressure projection
10:     $q_{v},q_{c},q_{p},\theta,\rho\leftarrow$ Thundercloud microphysics update Eqs.(1–15)
11:     if $S\cdot\rho>E_{tr}$ then $\triangleright$ Electrification process Eq.(16–18)
12:         $V,\rho_{\text{neut}}\leftarrow$ lightning discharge Eqs.(30–31),(19)
13:         $S\cdot\rho\leftarrow$ Apply lightning neutralization Eq.(20)
14:     end if
15:end procedure
The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1. Figure 3 visualizes the key interrelationships governing thunderstorm development. We first present our methodology for simulating the life cycle of a Mesoscale Convective System (MCS), followed by an explanation of the dynamics of lightning channel growth triggered when the electric field exceeds a defined threshold.
5.1. Mesoscale Convective System Cycle
The input fields for the MCS system, depicted in Figure 5, serve as the foundation for initializing state variables, including potential temperature ($\theta$), pressure ($p$), charge density ($\rho$), velocity field ($u$), and hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$).The solver updates these variables iteratively, producing an output that represents the evolved state of the MCS, incorporating changes driven by thundercloud microphysics and the electrification processes.
Figure 5. Different types of sources used as input for our MCS solver,temperature and vapor field are sampled from perlin noise patterns. (a) Height field derived from satellite data, (b) ground temperature distribution, (c) vapor field distribution.
The procedure begins by updating the atmospheric background conditions,through governing equations (Eqs. 22–23).
The velocity field is then advected and diffused.Thermal buoyancy is computed (Eq. 27) and integrated into the velocity field, alongside contributions from wind and vorticity confinement forces.
Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported according to the evolving velocity field through advection. Hydrometeor quantities ($q_{v}$, $q_{c}$, $q_{p}$) are transported through advection based on the evolving velocity field. For efficient pressure projection, we employ the compact Poisson filter scheme (Rabbani et al., 2022), a GPU-friendly method for solving large-scale sparse linear systems. Cloud microphysics are subsequently updated, incorporating processes such as condensation, evaporation, and precipitation, governed by parameterized equations (Eqs. 1–15).
Electrification is initiated when the charge density, modulated by the suppress factor ($S\cdot\rho$), exceeds the electric field threshold ($E_{tr}$, Eqs. 16–18). This triggers lightning discharges, which are detailed in the next subsection. Following lightning activity, the neutralized charge density ($\rho_{\text{neut}}$, Eqs. 30–31, 19) is updated, redistributing charge density ($S\cdot\rho$, Eq. 20). The suppress factor is recalibrated to reflect the diminished likelihood of subsequent lightning events as the system dissipates.
5.2. Lightning Channel Growth
The growth of the lightning channel begins at a trigger point, randomly selected from locations where the electric field exceeds a predefined threshold. Inspired by the DBM model proposed by (Kim and Lin, 2007), the lightning channel stops growing once it connects to a positively charged object. This approach not only enhances flexibility in controlling the final structure of the lightning channel but also improves computational efficiency. The positive charge is distributed along the top of the height field.
@@ 5.2. Lightning Channel Growth
The solution is computed on a volumetric grid, and the results are subsequently converted from grid points to geometric points. To create a more natural branching structure, a random dithering technique is applied to these points. The points are then connected to form the final geometry of the lightning branches.
@@ 6. VISUALIZATION
Our tool is implemented as a HDA(Houdini Digital Asset) using Houdini 20.0.625’s microsolver framework with OpenCL for GPU acceleration. The hardware setup includes NVIDIA® GeForce® RTX A6000 GPU,13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
The simulation results demonstrate various thunderstorm and lightning phenomena, integrated with scenarios inspired by real-world weather events. As shown in Table 1, we summarize the lightweight parameter set used for the scenes presented in this section. All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.
Table 1. Overview of the lightweight parameter set used in the scenes presented in this section. $D$: World dimensions (km); $R$: Volume resolution; $T_{\text{mcs}}$: Computation time in seconds per frame(s).
@@
6(a) Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
6(b) Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
6(c) Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
6(d) Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
7(a) Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
7(b) New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
7(c) Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
7(d) California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
6.1. Thunderstorms Variation
We simulate four common types of thunderstorms that occur within a mesoscale convective system (MCS), capturing their distinctive characteristics and dynamic behaviors:
@@ 6.1. Thunderstorms Variation
Figure 6. Comparison between simulation results (a-d) and realistic observations (e-h) for four storm types: single cell, multicell, squall line, and supercell. This layout highlights the structural similarities between our simulated storms and their real-world counterparts.
These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within an MCS. The visual comparison of these thunderstorm types is presented in Figure 6, highlighting their structural differences and unique features.
@@ 6.2. Severe Weather Phenomena
To explore the impact of thunderstorms in real-world scenarios, we reference severe weather events in geographically diverse locations. These events illustrate the variability and intensity of thunderstorms in different environments:
@@
Figure 7. Simulation cloudscapes inspired by real-world weather events. (a) Florida Biscayne National Park, (b) New Mexico Chiricahua Mountains, (c) Japan Fuji-Hakone-Izu National Park, (d) California coastal atmospheric river. These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.
  • Biscayne National Park, Florida (25.3692, -80.243): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
  • Chiricahua Mountains, New Mexico (35.7943, -106.443): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
  • Fuji-Hakone-Izu National Park, Japan (37.745, -119.54): On January 10, 2011, a sudden winter ;p‘thunderstorm formed in this region, resulting in brief but intense snowfall.
  • Monterey Bay, California (36.6197, -121.906): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather. The corresponding visualizations are presented in Figure 7.

### v4

@@ 4.2. Hydrometeor Electrification
Building on the Reynolds-Brook theory of thunderstorm electrification (Latham and Miller, 1965), the fair weather electric field induces opposing charges on precipitation particles. As these particles collide and coalesce, their velocities and directions influence partial charge neutralization, leaving a residual net charge.
This theory underpins the following mathematical model for electrification, with charge density $\rho$ defined as:
(13) $\rho=k\alpha(1-\alpha)\left(q_{c}^{2}\boldsymbol{u}_{c}-q_{p}^{2}\boldsymbol{u}_{p}\right)e^{-10q_{v}},$
@@
where $k$ is a user-defined constant, $\boldsymbol{u}_{c}$ and $\boldsymbol{u}_{p}$ are their respective velocities of cloud and precipitation. The threshold electric field $E_{tr}$ required for lightning initiation depends on the altitude $\rho_{A}(h)$ as defined by (Marshall et al., 1995):
(14) $E_{tr}=\pm 167\,\rho_{A}(h)\quad\text{where}\quad\rho_{A}(h)=1.208\exp\left(\frac{-h}{8.4}\right).$
To analyze the electrodynamics in the atmosphere, we solve the Poisson equation, a fundamental equation derived from Gauss’s law in electrostatics. This equation describes the simplified spatial variation of the potential $V$ due to charge density $\rho$(Kim and Lin, 2007):
(15) $\nabla^{2}V=-4\pi\rho.$
@@
Charge neutralization processes, occurring after a lightning discharge(Barthe and Pinty, 2007), are governed by the following relationship:
(16) $\delta\rho=\begin{cases}\pm\left(|\rho|-\rho_{\text{excess}}\right),&\text{if }|\rho|>\rho_{\text{excess}},\\ 0,&\text{if }|\rho|\leq\rho_{\text{excess}},\end{cases}$
@@
where $\delta\rho$ represents the net charge change, and $\rho_{\text{excess}}$ denotes the threshold for excess charge density. The total neutralized charge density, accounting for the collective contributions of lightning growth points, is expressed as:
(17) $\delta\rho_{\text{neut}}=\frac{1}{N}\left|(2\alpha-1)q_{c}q_{p}\right|\sum_{n}\delta\rho.$
@@
Finally, a suppression factor $S$ is introduced to modulate lightning activity as thunderstorms dissipate:
(18) $S=\frac{1}{1+k\cdot|\delta\rho_{\text{neut}}|}.$
@@ 4.3. Atmospheric Background
(19) $\gamma_{\text{th}}=Y_{\text{vapor}}\gamma_{\text{vapor}}+\left(1-Y_{\text{vapor}}\right)\gamma_{\text{air}},$
@@
(20) $T(h)=\begin{cases}T_{0}+\Gamma_{0}h,&0\leq h\leq h_{1},\\ T_{0}+\Gamma_{0}h_{1}+\Gamma_{1}(h-h_{1}),&h_{1}\leq h,\end{cases}$
@@
(21) $p(h)=p_{0}\left(1-\frac{\Gamma h}{T_{0}}\right)^{\frac{g}{RT_{0}}},$
@@
(24) $T_{\mathrm{th}}(h)=T_{0}\left(\frac{p(h)}{p_{0}}\right)^{\frac{\gamma_{\mathrm{th}}-1}{\gamma_{\mathrm{th}}}}.$
@@ 4.4. Fluid Dynamics
The motion of atmospheric fluids is governed by the incompressible Navier-Stokes equations, which represent the principles of momentum and mass conservation in a fluid medium (Wendt, 2008). These equations account for the effects of inertial forces, pressure gradients, viscosity, and external forces:
(26) $\frac{\partial(\varrho\boldsymbol{u})}{\partial t}+\nabla\cdot(\varrho\boldsymbol{u}\boldsymbol{u})=-\nabla p+\varrho\left(\mu\nabla^{2}\boldsymbol{u}+\boldsymbol{f}\right),\quad\nabla\cdot\boldsymbol{u}=0,$
where $\varrho$ represents the fluid density, $\boldsymbol{u}$ is the velocity, $p$ is the pressure, $\mu$ denotes the dynamic viscosity, and $\mathbf{f}$ encompasses buoyancy and other external forces. The equation $\nabla\cdot\boldsymbol{u}=0$ represents the continuity equation, ensuring the conservation of mass in incompressible flows.The pressure term, $p$, is determined via the Helmholtz–Hodge decomposition by solving the Poisson equation:
(27) $\nabla^{2}p=\nabla\cdot\boldsymbol{u}.$
@@ 5. ALGORITHMICS
The theoretical framework discussed in the previous section is translated into a numerical procedure, as outlined in Algorithm 1. Figure 2 visualizes the key interrelationships governing MCS development. We first present our algorithm for simulating the life cycle of a MCS, followed by an explanation of the implementation details of our procedure.
5.1. Mesoscale Convective System Cycle
The evolution of the MCS system follows Algorithm 1, iteratively updating state variables, including potential temperature $\theta$, pressure $p$, charge density $\rho$, velocity field $\boldsymbol{u}$, and hydrometeor quantities $q_{v}$, $q_{c}$, and $q_{p}$. The process begins with updating atmospheric background conditions $\theta,p,\rho$ following Eqs. (20–21). The velocity field $\boldsymbol{u}$ is then advected and diffused, followed by the computation of thermal buoyancy $b$ based on Eq. (25). Buoyancy, wind, and vorticity confinement forces are subsequently integrated into velocity field $\boldsymbol{u}$ to reflect key atmospheric dynamics.
Algorithm 1 Thunderscapes Algorithm
1:Input: Current MCS state $(\theta,p,\rho,\boldsymbol{u},q_{v},q_{c},q_{p})$.
@@ 5.1. Mesoscale Convective System Cycle
4:     $\theta,p,\rho\leftarrow$ Update atmospheric background conditions Eqs.(20–21)
5:     $\boldsymbol{u}\leftarrow$ Advect and diffuse velocity field
6:     $\boldsymbol{b}\leftarrow$ Compute thermal buoyancy Eq.(25)
7:     $\boldsymbol{u}\leftarrow\boldsymbol{u}+\boldsymbol{b}+\boldsymbol{f}_{w}+\boldsymbol{f}_{v}$triangleright Apply buoyancy, wind, vorticity confinement
@@
9:     $\boldsymbol{u}\leftarrow$ Pressure projection Eq.(27)
10:     $q_{v},q_{c},q_{p},\theta\leftarrow$ Update cloud microphysics and temperature Eqs.(2–5)
11:     $\rho\leftarrow$ Hydrometeor electrification Eqs.(13)
12:     if $S\cdot\rho>E_{tr}$ then Eq.(14)
13:         $V,\rho_{\text{neut}}\leftarrow$ lightning discharge Eqs.(15–17)
14:         $S\leftarrow$ Apply lightning neutralization Eq.(18)
15:     end if
16:end procedure
Hydrometeor quantities $q_{v},q_{c},q_{p}$ are transported through advection in accordance with the updated velocity field $\boldsymbol{u}$. To ensure mass conservation, the velocity field $\boldsymbol{u}$ is corrected via pressure projection following Eq. (27). Next, cloud microphysics are updated by solving parameterized equations based on Eqs. (2–5) to account for condensation, evaporation, and precipitation.
Hydrometeor electrification is modeled using parameterized equations following Eq. (13). When the charge density, modulated by the suppress factor $S\cdot\rho$, exceeds the electric field threshold $E_{tr}$ based on Eq. (14), lightning discharge is triggered. During the discharge, lightning channel growth is stochastically modeled using principles from the Dielectric Breakdown Model (DBM) (Kim and Lin, 2007). Finally, neutralized charge density $\rho_{\text{neut}}$ is updated following Eqs. (15–17), and the suppress factor is recalibrated to reduce the likelihood of subsequent lightning events as thunderstorms dissipate based on Eq. (18).
5.2. Implementation
Our tool is implemented as a Houdini Digital Asset (HDA) using Houdini 20.0.625’s microsolver framework, with OpenCL for GPU acceleration. The hardware setup comprises an NVIDIA® GeForce® RTX A6000 GPU, a 13th Gen Intel® Core™ i9-13900 processor, and 128GB of RAM.
The framework is designed to align with the traditional workflow of digital artists, utilizing custom volume fields as input sources, including temperature field, vapor field, and height field.These fields, representing ground properties, serve as the foundation for driving the MCS simulation. Our framework operates on a uniform 3D grid, utilizing a semi-Lagrangian scheme to advect system quantities. To simulate horizontal thunderstorm movement, we apply a uniform horizontal wind field. Combined with vertical buoyancy forces, this approach produces a more realistic depiction of thunderstorm surges.
To mitigate numerical dissipation, which can cause small-scale vortices to vanish prematurely, we integrate vorticity confinement techniques (Miyazaki et al., 2002). This enhancement ensures that fine-scale details, such as swirling motions within the storm, are preserved throughout the simulation.
@@ 5.2. Implementation
Lightning channel growth during hydrometeor discharge events is modeled using the Dielectric Breakdown Model (DBM) proposed by Kim et al. (Kim and Lin, 2007). The resulting lightning channels are converted from grid points to geometric points. To achieve a more natural branching structure, a random dithering technique is applied to these points, which are subsequently connected to form the final geometry of the lightning branches.
Both the sparse Poisson problem for fluid pressure projection and the DBM processes are efficiently handled using the compact Poisson filter (Rabbani et al., 2022), a GPU-friendly method specifically optimized for solving large sparse linear systems.
@@ 6. VISUALIZATION
The simulation results demonstrate various thunderstorm phenomena, integrated with scenarios inspired by real-world weather events. As shown in Table 1, we summarize the spatial scale used for the scenes presented in this section. All renders are produced using Houdini’s native volumetric rendering engine to ensure high-quality visual outputs.For dynamic details, please refer to our supplementary video.
Table 1. The table provides an overview of the spatial scale used in the scenes presented in this paper. Moreover, resolution ($R$) and computation time ($T_{\text{mcs}}$) measured in seconds per frame are listed. A constant time step size of $\Delta t=1\,\mathrm{min}$ is used. The background parameters defining the atmosphere are set to $\Gamma=-8.5\,\mathrm{K}/\mathrm{km}$ and $h_{1}=8\,\mathrm{km}$.
@@
 3 Single Cell $18\times 12\times 18$ $450\times 301\times 450$ 0.48
 3 Multicell $20\times 15\times 20$ $500\times 376\times 500$ 0.81
 3 Squall Line $30\times 15\times 30$ $500\times 251\times 500$ 0.54
 3 Super Cell $30\times 15\times 30$ $500\times 251\times 500$ 0.68
 4 Florida $11\times 12\times 11$ $287\times 309\times 287$ 0.24
 4 New Mexico $11\times 13\times 11$ $383\times 438\times 383$ 0.59
 4 Japan $11\times 13\times 11$ $383\times 444\times 383$ 0.64
 4 California $10\times 7\times 10$ $340\times 247\times 340$ 0.28
6.1. Thunderstorm Variation
We simulate four common types of thunderstorms that occur within a MCS according to the National Severe Storms Laboratory (NSSL) [^1] https://www.nssl.noaa.gov/education/svrwx101/thunderstorms/types/, capturing their distinctive characteristics and dynamic behaviors:
@@ 6.1. Thunderstorm Variation
Refer to caption(i)
Refer to caption(j)
Refer to caption(k)
Refer to caption(l)
Figure 3. Comparison between simulation results (a-d) and realistic observations (e-h) for four storm types: single cell, multicell, squall line, and supercell. This layout highlights the structural similarities between our simulated storms and their real-world counterparts. Additionally, the last row (i-l) illustrates the ground properties (including the temperature field and vapor field initialized with a consistent pattern based on a flat heightfield) associated with each thunderstorm type.
These simulated variations allow us to explore the diverse behaviors and impacts of thunderstorms within a MCS. Figure 3 presents a visual comparison along with the corresponding parameter sets, showcasing the structural differences and distinctive features of these thunderstorm types.
@@ 6.2. Severe Weather Phenomena
To explore the impact of MCSs in real-world scenarios, we reference severe weather events from geographically diverse locations using the Storm Events Database provided by the National Oceanic and Atmospheric Administration (NOAA) [^2] https://www.ncdc.noaa.gov/stormevents. These events highlight the variability and intensity of thunderstorms across different environmental conditions and seasons:
  • Biscayne National Park, Florida (25.3692°N, -80.243°W): On March 13, 1993, a tropical storm occurred in this region during the spring, bringing heavy rainfall and strong winds.
  • Chiricahua Mountains, New Mexico (35.7943°N, -106.443°W): On July 14, 1989, a severe thunderstorm swept through this mountainous area during the summer, producing heavy rain, hail, and strong winds.
  • Fuji-Hakone-Izu National Park, Japan (37.745°N, -119.54°W): On January 10, 2011, a sudden winter thunderstorm formed in this region, resulting in brief but intense snowfall.
  • Monterey Bay, California (36.6197°N, -121.906°W): On December 2, 2012, an atmospheric river brought heavy rainfall and thunderstorms along the California coast, accompanied by strong gusty winds.
@@
Refer to caption(e)
Refer to caption(f)
Refer to caption(g)
Refer to caption(h)
Figure 4. Simulation inspired by real-world weather events:(a) Florida’s Biscayne National Park, (b) New Mexico’s Chiricahua Mountains, (c) Japan’s Fuji-Hakone-Izu National Park, (d) California’s Monterey Bay. These visualizations highlight the geographical diversity and meteorological phenomena captured in our simulation framework.Additionally, the last row (e-h) illustrates the ground properties (including the temperature field and vapor field, both initialized with a consistent pattern derived from a specific regional heightfield) associated with each region.
These events serve as the basis for our simulations, showcasing the capability of our model to replicate the diverse and complex phenomena associated with severe weather. The corresponding visualizations and parameter sets are presented in Figure 4.